Question Three: In 2016, the Westgate Construction Company entered into a contract to construct a road for Santa Clara Countyfor $10,000,000. The road was completed in 201 contract is as follows: 8. Information related to the 2018 2016 2,400,000 3,600,000 2,200,00 5,600,000 2,000,000 4,000,000 4,000,000 ,800,000 3,600,000 4,600,000 2017 Cost incurred during the year Estimated costs to complete as of year-end Billings during the year Cash collections during the year 2,000,000 Westgate recognizes revenue over time according to percentage of completion. 1. briefly explain the difference between recognizing revenue over time and upon project completion when accounting for long-term construction contracts.
Answer #1

Answer

Requirement 1

  • When the revenue is recognised over the time of the project, each year, part of total value of project is considered as a Revenue based on ‘percentage’ of completion.
  • When revenue is recognised on completion, no revenue is recognised until and unless whole project gets complete.
  • The total Gross Profit (Loss) on project stays the same over the life of project.
  • Considering the data in the question, following workings will demonstrate the differences in Percentage of Completion Method and Completion method.

---Percent of completion method

Working

Year end 2016

Year end 2017

Year end 2018

A

Contract Price

$     10,000,000.00

$    10,000,000.00

$        10,000,000.00

B

Cost Incurred to Date

$       2,400,000.00

$       6,000,000.00

$          8,200,000.00

C

Estimated cost yet to be incurred to complete the contract

$       5,600,000.00

$       2,000,000.00

$                                -  

D = B+C

Total Cost

$       8,000,000.00

$       8,000,000.00

$          8,200,000.00

E = (B/D) x 100

% of Completion

30.00%

75.00%

100.00%

F = A x E

Revenue to date

$       3,000,000.00

$       7,500,000.00

$        10,000,000.00

G

Revenue of Previous year

$                            -  

$       3,000,000.00

$          7,500,000.00

H = F - G

Net Revenue this year

$       3,000,000.00

$       4,500,000.00

$          2,500,000.00

I (=B)

Cost to date

$       2,400,000.00

$       6,000,000.00

$          8,200,000.00

J

Cost to date of previous year

$                            -  

$       2,400,000.00

$          6,000,000.00

K = I - J

Net Cost for the year

$       2,400,000.00

$       3,600,000.00

$          2,200,000.00

L = H - K

Gross Profits

$           600,000.00

$          900,000.00

$              300,000.00

---Completion Method

Contract Price

$     10,000,000.00

Less: Total Cost

$       8,200,000.00

Gross Profit on Contract to be recorded in 2018

$       1,800,000.00

Notes:

1) No profit or loss will be recorded in the year 2017 and 2018 under Compete Contract Method

2) The Profits or loss under both methods will be Exactly same, The only diffrence is Profit or loss is recognised in year of completion in case of Completed Contract method and under Percentage on completion method Profit or loss is recognised in all years in which work is done.